# Standing-Wave Dynamics in Low-Frequency Breathing of a Turbulent Separation Bubble

### v2

@@ Abstract
This study investigates the low-frequency dynamics of a turbulent separation bubble (TSB) forming over a backward-facing ramp, with a focus on large-scale coherent structures associated with the so-called ’breathing motion’. Using time-resolved particle image velocimetry (PIV) in both streamwise and spanwise planes, we examine the role of sidewall confinement, an aspect largely overlooked in previous research. Spectral proper orthogonal decomposition (SPOD) of the streamwise velocity field reveals a dominant low-rank mode at low Strouhal numbers ($St<0.05$), consistent with prior observations of TSB breathing. Strikingly, the spanwise-oriented PIV data uncover a previously unreported standing wave pattern, characterised by discrete spanwise wavenumbers and nodal/antinodal structures, suggesting the presence of spanwise resonance. To explain these observations, we construct a resolvent-based model that imposes free-slip conditions at the sidewall locations by superposing left- and right-traveling three-dimensional modes. The model accurately reproduces the spanwise structure and frequency content of the measured SPOD modes, demonstrating that sidewall reflections lead to the formation of standing wave-like patterns. To gain further insight into the driving mechanisms of the low-frequency dynamics, a global stability analysis is performed, revealing a zero-frequency eigenmode whose growth rate depends on the spanwise wavenumber. This eigenmode originates from a centrifugal instability and gives rise to the observed low-frequency breathing. Downstream, the associated coherent structures are further amplified through non-modal lift-up mechanisms. Our findings highlight the critical influence of spanwise boundary conditions on the selection and structure of low-frequency modes in TSBs. This has direct implications for both experimental and numerical studies, particularly those relying on spanwise-periodic boundary conditions, and offers a low-order framework for predicting sidewall-induced modal dynamics in separated flows.
@@ 1 Introduction
TSBs can be classified into two main categories: geometry-induced and adverse pressure gradient (APG)-induced separation bubbles. Geometry-induced TSBs arise at sharp geometric discontinuities such as backward-facing steps (EatonJohnston1982) and rectangular leading-edge configurations (CherryHillier_1984_JFM; Kiya_Sasaki_1983). APG-induced TSBs occur in configurations such as airfoils (Delery1985_shock), backward-facing ramps (Kaltenbach1999StudySimulation; WeissAIAAJ2022), or flat plates where an adverse pressure gradient drives separation (Patrick1987_Nasa; NaMoin1998JFM; Wu2019; Cura2024; Abe2017; MohammedTaifur_Weiss_2016). One specific instance of an APG-induced TSB is the trailing-edge separation observed on airfoils, where the flow detaches in the rear part of the airfoil and reattaches near the trailing edge. Low-frequency oscillations of these trailing-edge TSBs were first noted by ZamanMcKinizieRumsey1989JfM and have since remained a topic of ongoing research (WandGhaemi2022_JfM; SarrasJfM2024).
@@
The mechanism driving the breathing dynamics is still an open question. Wu2019 used a direct numerical simulation (DNS) to study a flat-plate TSB with two setups: one using suction-only separation and another using a suction-and-blowing profile to force reattachment. Dynamic mode decomposition (DMD) revealed streamwise-elongated structures potentially linked to Görtler instability, but their observed breathing frequencies ($St\approx 0.45$) were significantly higher than those reported in other studies. More recently, Cura2024 performed global stability (LSA) and resolvent analyses (RA) on a DNS dataset from Coleman2018NumericalNumber. Their global stability analysis found no unstable eigenmodes, but the least damped mode occurred at zero frequency with small, nonzero spanwise wavenumbers. Resolvent analysis showed strong agreement with experimental databases, suggesting that forced dynamics are influenced by a weakly damped global mode.
SarrasJfM2024 analyse stall cells using global linear stability analysis on a RANS mean flow and compare their findings with wind tunnel experiments. Their findings reveal stationary three-dimensional global modes becoming unstable only within certain ranges of angles of attack and particular spanwise wavenumbers.
Bugeat2022JfM investigated the low-frequency breathing dynamics of a laminar oblique SBLI and conducted a RA on a steady 2D base flow. They found a modal mechanism for large spanwise wavelengths of the order of the separation bubble length with large gain values. At smaller spanwise wavelengths, in the order of the boundary layer thickness, they identified the growth of streaks, which they attributed to a non-modal mechanism. Marquet2009DirectNon examined a laminar separation bubble behind a smooth ramp and found that destabilization occurred due to a stationary three-dimensional mode with a spanwise wavenumber. This mode was linked to the formation of high- and low-speed streaks, which they identified as a lift-up effect.
@@
Gallaire2007 similarly analysed a laminar separation bubble developing over a bump at $Re=400$ and identified stationary three-dimensional unstable eigenmodes. Their study further examined whether these modes were associated with a centrifugal-type instability.
@@
SavarinoJFM2025 investigated a transitional flat-plate separation-bubble flow using both linear stability and resolvent analyses. They identified zero-frequency eigenmodes consistent with a centrifugal mechanism, while the corresponding resolvent spectrum exhibited a secondary hump at higher spanwise wavenumbers, that is related to a non-modal lift-up process. Additionally, they reported a configuration in which both centrifugal and lift-up mechanisms appeared to be active simultaneously.
Although Bugeat2022JfM, Marquet2009DirectNon, SavarinoJFM2025, Gallaire2007, SarrasJfM2024, and Cura2024 investigate different flow configurations, their findings consistently indicate that the amplification behaviour of the low-frequency breathing mechanism in separation bubbles exhibits a clear spanwise dependency.
@@
Given the strong influence of spanwise wavenumber on amplification and stability mechanisms, the spanwise dimension and boundary conditions likely play a crucial role. While the impact of sidewalls and the distinction between periodic and solid-wall boundary conditions on low-frequency modes have been explored in SBLI research (Rabey2019JfM; Garnier2009), where they were attributed to corner effects, their influence remains largely unexamined in subsonic TSB studies. Many numerical studies (Cura2024; Wu2019) utilise periodic sidewalls, as they minimise spurious boundary effects, reduce computational costs, and preserve statistical homogeneity. However, periodic boundary conditions impose constraints by allowing only integer spanwise wavenumbers. Experimental setups, in contrast, typically have solid sidewalls, that influence the flow dynamics.
@@ 2.1 Experimental setup
The experiments were conducted in a low-speed closed-loop wind tunnel. Thanks to its temperature control, the Reynolds number was kept constant throughout the measurements at $Re_{\theta}\approx 1000$, based on the inflow velocity $u_{\infty}=20\,\mathrm{m/s}$ and the momentum thickness $\theta\approx 0.8\,\mathrm{mm}$ that was measured using hot-wire anemometry in the test section centre line, $200\,\mathrm{mm}$ upstream of the ramp. The test section and the coordinate system are shown in Figure 1. Upstream of the ramp, the channel has a height of $h=400\,\mathrm{mm}$. The entire domain has a spanwise width of $L=600\,\mathrm{mm}$. Flow separation inside the test section was induced by an adverse pressure gradient due to the linear widening of the cross-section shown in Figure 1. Reattachment occurs downstream of the backward-facing ramp, which was tilted by an angle of $\alpha=$ 20 degrees, as the separated shear-layer impinges onto the horizontal plate. The average streamwise length of the separation bubble on the centre line is $L_{b}=200\,\mathrm{mm}$. It is taken as the characteristic length to scale the frequency (using $St=fL_{b}/u_{\infty}$) and spanwise wavenumbers $\beta$ of the coherent structures. A key focus of this study is to examine how sidewalls and spanwise domain length influence mode selection and the geometric scaling of coherent structures. Therefore, the spanwise domain length $L$ is additionally referenced in many figures.
@@
It was recently shown by (SteinfurthPINN_PoF2024) that the mean flow topology for this setup is three-dimensional, which is prominently manifested in a U-shaped mean separation line as seen in Figure 2(a). In SteinfurthPINN_PoF2024, a physics-informed neural network (PINN) was trained on pressure, wall-shear stress, and three-component velocity data obtained from PIV to reconstruct the 3D mean flow. Figure 2(a) is based on this model and illustrates that flow separation occurs further upstream near the sidewalls, resulting in a longer separated region in these areas and a shorter separated region near the symmetry plane. The spanwise inhomogeneity of the mean flow is highlighted in Figure 2(b), while the combination of wall shear stress data and oil-film images in Figure 2(c) further emphasises the three-dimensional nature of the flow.
@@ 2.2 Data processing using spectral proper orthogonal decomposition
In this study, we apply the SPOD algorithm developed by Lumley1970; SchmidtColonius2020. For the symmetry plane, both $u^{\prime}(\boldsymbol{x},t)$ and $v^{\prime}(\boldsymbol{x},t)$ velocity components are used as input. The time series is segmented into overlapping blocks of 256 snapshots with 50% overlap. Each segment is windowed to reduce spectral leakage, and a Fourier transform is applied. The cross-spectral density (CSD) matrices are computed for each frequency based on the Fourier mode estimates of all blocks. Finally, an eigendecomposition of the CSD matrix yields the SPOD modes ranked by energy, analogous to standard POD but with spectral resolution as the CSD matrix is formed for each frequency separately.
@@ 3.1 RANS Setup
The initial condition for the turbulence kinetic energy $k$ and the turbulence specific dissipation rate $\omega$ can be set based on characteristic length and velocity scales and a turbulence intensity value (openFoamUserGuide_kOmegaSST). The latter describes the ratio between the root mean square of the turbulent velocity fluctuations and the mean velocity in percent. A value of 3.75% was carefully chosen aiming to minimize the difference in separation and reattachment points and separation bubble size between the RANS solution and the PIV mean field. This approach takes advantage of the strong influence of freestream turbulence on the separation point of the boundary layer from the surface. In particular, higher values of free stream turbulence cause a larger momentum transfer normal to the streamwise direction of the flow. This results in a longer capability of the flow to stay attached. So the separation point is moved further downstream, resulting in a smaller separation bubble. Lower freestream turbulence values move the separation point upstream, resulting in a larger separation bubble.
@@
Figure 4 shows the $\bar{u}$- and $\bar{v}$-components of the resulting RANS mean field alongside a comparison with the PIV mean field. Slight deviations can be observed in the recirculation region: the PIV data show a slightly larger separation bubble in the crosswise direction with slightly stronger backflow and the RANS mean field displays a thin separation bubble in the upstream part which is not visible in the PIV mean field. This latter observation may be attributed to the presence of spurious measurement data near the wall in the current dataset, caused by low signal-to-noise ratios due to strong light reflections. This experimental shortcoming does not negatively affect the SPOD analysis, since the breathing mode unfolds further away from the wall, as will be shown later. Iso-contours of $\bar{u}/u_{\infty}=0.2$ and $\bar{v}/u_{\infty}=-0.05$ are included to allow for a more quantitative comparison. While the backflow region is smaller in the RANS field, the selected contour levels indicate closer agreement between the two datasets. With the chosen RANS model using the $k$-$\omega$-SST turbulence model and adaption of the turbulence intensity value at the inlet of the domain, it was not possible to better approximate the crosswise extent of the bubble without significantly altering the separation and reattachment points. Similar challenges in matching RANS-predicted separation bubble behaviour with experimental observations have been reported for smooth-body separation, where accurate prediction remains difficult and continues to be an active topic of research (williams_Experimental_2020; gray_Experimental_2023; Klopsch2025). SarrasJfM2024 observed comparable discrepancies and showed that improved agreement could be achieved by applying adjoint-based data assimilation with flexibility in the turbulence model.
@@ 3.2 Global Linear Stability Analysis
To investigate modal amplification mechanisms in the flow, a global linear stability analysis (LSA) of the turbulent mean flow is performed. This approach examines the time-asymptotic dynamics of harmonic perturbations that evolve around the temporal mean state by performing an eigenvalue decomposition of the linearized Navier–Stokes operator, which yields growth rates and corresponding mode shapes. Linear stability analysis has become a common framework to study coherent flow dynamics (Sipp2010; Barkley2006) and has recently been applied to separation bubble flows (Cura2024; SavarinoJFM2025; SarrasJfM2024).

### v3

@@ Abstract
This study investigates the low-frequency dynamics of a turbulent separation bubble (TSB) forming over a backward-facing ramp, with a focus on large-scale coherent structures associated with the so-called ’breathing motion’. Using time-resolved particle image velocimetry (PIV) in both streamwise and spanwise planes, we examine the role of sidewall confinement, an aspect largely overlooked in previous research. Spectral proper orthogonal decomposition (SPOD) of the streamwise velocity field reveals a dominant low-rank mode at low Strouhal numbers ($St<0.05$), consistent with prior observations of TSB breathing. Strikingly, the spanwise-oriented PIV data uncover a previously unreported standing wave pattern, characterised by discrete spanwise wavenumbers and nodal/antinodal structures, suggesting the presence of spanwise resonance. To explain these observations, we construct a resolvent-based model that imposes free-slip conditions at the sidewall locations by superposing left- and right-traveling three-dimensional modes. The model accurately reproduces the spanwise structure and frequency content of the measured SPOD modes, demonstrating that sidewall reflections lead to the formation of standing wave-like patterns. Global stability analysis reveals a zero-frequency eigenmode originating from a centrifugal instability, giving rise to the observed low-frequency breathing. Downstream, the associated coherent structures are further amplified through non-modal lift-up mechanisms. Our findings highlight the critical influence of spanwise boundary conditions on the selection and structure of low-frequency modes in TSBs. This has direct implications for both experimental and numerical studies relying on spanwise-periodic boundary conditions and offers a low-order framework for predicting sidewall-induced modal dynamics in separated flows.
@@ 1 Introduction
TSBs can be classified into two main categories: geometry-induced and adverse pressure gradient (APG)-induced separation bubbles. Geometry-induced TSBs arise at sharp geometric discontinuities such as backward-facing steps (EatonJohnston1982) and rectangular leading-edge configurations (CherryHillier_1984_JFM; Kiya_Sasaki_1983). APG-induced TSBs occur in configurations such as airfoils (Delery1985_shock), backward-facing ramps (Kaltenbach1999StudySimulation; WeissAIAAJ2022), or flat plates where an adverse pressure gradient drives separation (Patrick1987_Nasa; NaMoin1998JFM; Wu2019; Cura2024; Cura2025; Abe2017; MohammedTaifur_Weiss_2016). One specific instance of an APG-induced TSB is the trailing-edge separation observed on airfoils, where the flow detaches in the rear part of the airfoil and reattaches near the trailing edge. Low-frequency oscillations of these trailing-edge TSBs were first noted by ZamanMcKinizieRumsey1989JfM and have since remained a topic of ongoing research (WandGhaemi2022_JfM; SarrasJfM2024).
@@
The mechanism driving the breathing dynamics is still an open question. Wu2019 used a direct numerical simulation (DNS) to study a flat-plate TSB with two setups: one using suction-only separation and another using a suction-and-blowing profile to force reattachment. Dynamic mode decomposition (DMD) revealed streamwise-elongated structures potentially linked to Görtler instability, but their observed breathing frequencies ($St\approx 0.45$) were significantly higher than those reported in other studies. Cura2024 performed global stability (LSA) and resolvent analyses (RA) on a DNS dataset from Coleman2018NumericalNumber. Their global stability analysis found no unstable eigenmodes, but the least damped mode occurred at zero frequency with small, nonzero spanwise wavenumbers. Resolvent analysis showed strong agreement with experimental databases, suggesting that forced dynamics are influenced by a weakly damped global mode. More recently, Cura2025 studied a family of turbulent separation bubbles of increasing size and showed that the growth rates scale with bubble dimensions, quantified by the reverse-flow magnitude and separation length. Unstable global modes were identified for the largest TSB extent. Another example in which a stationary three-dimensional mode has been identified in TSB research is provided by studies of stall cells. SarrasJfM2024 analyse stall cells using global linear stability analysis on a RANS mean flow and compare their findings with wind tunnel experiments. Their findings reveal stationary three-dimensional global modes becoming unstable only within certain ranges of angles of attack and particular spanwise wavenumbers.
The role of stationary three-dimensional eigenmodes as a driving feature of separation-bubble dynamics has already been demonstrated for laminar separation bubbles. Theofilis2000 analysed a laminar flat-plate flow and reported unstable, stationary, spanwise-periodic three-dimensional eigenmodes. Rodrguez2010 identified the same global mode and showed that it induces a spanwise modulation of the flow. Subsequently, Rodrguez2013 attributed this behaviour to a three-dimensional centrifugal instability and demonstrated that the associated growth rates depend on the reverse-flow intensity and the separation-bubble length.
@@
Gallaire2007 similarly analysed a laminar separation bubble developing over a bump at $Re=400$ and identified stationary three-dimensional unstable eigenmodes. Their study further examined whether these modes were associated with a centrifugal-type instability. Marquet2009DirectNon examined a laminar separation bubble behind a smooth ramp and found that destabilization occurred due to a stationary three-dimensional mode with a spanwise wavenumber. This mode was linked to the formation of high- and low-speed streaks, which they identified as a lift-up effect.
@@
In the transitional regime, SavarinoJFM2025 investigated a flat-plate separation-bubble flow using both linear stability and resolvent analyses. They identified zero-frequency eigenmodes consistent with a centrifugal mechanism, while the corresponding resolvent spectrum exhibited a secondary hump at higher spanwise wavenumbers, that is related to a non-modal lift-up process. Additionally, they reported a configuration in which both centrifugal and lift-up mechanisms appeared to be active simultaneously.
Bugeat2022JfM investigated the low-frequency breathing dynamics of a laminar oblique SBLI and conducted a RA on a steady 2D base flow. They found a modal mechanism for large spanwise wavelengths of the order of the separation bubble length with large gain values. At smaller spanwise wavelengths, in the order of the boundary layer thickness, they identified the growth of streaks, which they attributed to a non-modal mechanism.
@@
Although Bugeat2022JfM, Marquet2009DirectNon, SavarinoJFM2025, Gallaire2007, SarrasJfM2024, and Cura2024 investigate different flow configurations, their findings consistently indicate that the amplification behaviour of the low-frequency breathing mechanism in separation bubbles exhibits a clear spanwise dependency. Given the strong influence of spanwise wavenumber on amplification and stability mechanisms, the spanwise dimension and boundary conditions likely play a crucial role. The impact of sidewalls and the distinction between periodic and solid-wall boundary conditions on low-frequency modes have been explored in SBLI research (Rabey2019JfM; Garnier2009), where they were attributed to corner effects. Similarly, in transitional backward-facing step configurations, Kaiktsis1991 and Barkley2002 reported that the type (solid vs. periodic) and spacing of sidewalls influences the spanwise modulation of the separation length. However, such influence remains largely unexamined in subsonic TSB studies. Many numerical studies (Cura2024; Wu2019) utilise periodic sidewalls, as they minimise spurious boundary effects, reduce computational costs, and preserve statistical homogeneity. However, periodic boundary conditions impose constraints by allowing only integer spanwise wavenumbers. Experimental setups, in contrast, typically have solid sidewalls, that influence the flow dynamics.
@@ 2.1 Experimental setup
The experiments were conducted in a low-speed closed-loop wind tunnel which maintains a constant Reynolds number throughout the measurements through temperature control. The test section and the coordinate system are shown in Figure 1. The free stream velocity of the test section is set to $u_{\infty}=20\,\mathrm{m/s}$. The momentum thickness $\vartheta\approx 0.7\,\mathrm{mm}$ is determined using hot-wire anemometry at $x=-200\,\mathrm{mm}$, yielding a Reynolds number of $Re_{\vartheta}\approx 1\ 000$. Upstream of the ramp, the channel has a height of $h=400\,\mathrm{mm}$. The domain has a spanwise width of $L=600\,\mathrm{mm}$. Flow separation inside the test section was induced by an adverse pressure gradient due to the linear widening of the cross-section shown in Figure 1. Reattachment occurs downstream of the backward-facing ramp, which was tilted by an angle of $\alpha=$ 20 degrees, as the separated shear-layer impinges onto the horizontal plate. The average streamwise length of the separation bubble on the centre line is approximately $L_{b}=200\,\mathrm{mm}$. This characteristic length scale is used to normalise the frequency (using $St=fL_{b}/u_{\infty}$) and spanwise wavenumbers $\beta$ of the coherent structures throughout this article. Since the key focus of this study is to examine how sidewalls and spanwise domain length influence mode selection and the geometric scaling of coherent structures, the spanwise domain length $L$ is additionally referenced in many figures.
@@
It was recently shown by SteinfurthPINN_PoF2024 that the mean flow topology for this setup is three-dimensional, which is prominently manifested in a U-shaped mean separation line as seen in Figure 2(a). They trained a physics-informed neural network (PINN) on pressure, wall-shear stress, and three-component velocity data obtained from PIV to reconstruct the 3D mean flow. Figure 2(a) is based on this model and illustrates that flow separation occurs further upstream near the sidewalls, resulting in a longer separated region in these areas and a shorter separated region near the symmetry plane. The spanwise inhomogeneity of the mean flow is highlighted in Figure 2(b), while the combination of wall shear stress data and oil-film images in Figure 2(c) further emphasises the three-dimensional nature of the flow.
@@ 2.2 Data processing using spectral proper orthogonal decomposition
In this study, we apply the SPOD algorithm developed by Lumley1970; SchmidtColonius2020. For the symmetry plane, both $u^{\prime}(\boldsymbol{x},t)$ and $v^{\prime}(\boldsymbol{x},t)$ velocity components are used as input. The time series is segmented into overlapping blocks of 200 snapshots with 50% overlap. Each segment is windowed to reduce spectral leakage, and a Fourier transform is applied. The cross-spectral density (CSD) matrices are computed for each frequency based on the Fourier mode estimates of all blocks. Finally, an eigendecomposition of the CSD matrix yields the SPOD modes ranked by energy, analogous to standard POD but with spectral resolution as the CSD matrix is formed for each frequency separately.
@@ 3.1 RANS Setup
The initial condition for the turbulence kinetic energy $k$ and the turbulence specific dissipation rate $\omega$ can be set based on characteristic length and velocity scales and a turbulence intensity value (openFoamUserGuide_kOmegaSST). The latter describes the ratio between the root mean square of the turbulent velocity fluctuations and the mean velocity in percent. A value of 3.75% was carefully chosen aiming to minimize the difference in separation bubble size between the RANS solution and the PIV mean field. This approach takes advantage of the strong influence of freestream turbulence on the separation point of the boundary layer from the surface. In particular, higher values of free stream turbulence cause a larger momentum transfer normal to the streamwise direction of the flow. This results in a longer capability of the flow to stay attached. So the separation point is moved further downstream, resulting in a smaller separation bubble. Lower freestream turbulence values move the separation point upstream, resulting in a larger separation bubble. This approach yields a similar inflow momentum thickness as in the PIV measurements. While at $x=-200\,\mathrm{mm}$, the RANS simulation indicates a larger value ($\vartheta\approx 1.1\,\mathrm{mm}$) compared to the experiment ($\vartheta\approx 0.7\,\mathrm{mm}$), the deviation is marginal at $x=-50\,\mathrm{mm}$ where $\vartheta\approx 0.9\,\mathrm{mm}$ (PIV) versus $\vartheta\approx 1.0\,\mathrm{mm}$ (RANS). Figure 4 shows the $\bar{u}$- and $\bar{v}$-components of the resulting RANS mean field alongside the PIV mean field. Iso-contours of $\bar{u}/u_{\infty}=0.2$ and $\bar{v}/u_{\infty}=-0.05$ are included to allow for a more quantitative comparison. While the flow topologies agree reasonably well, slight deviations can be observed in the recirculation region: the PIV data show a slightly larger separation bubble in the crosswise direction with stronger reverse flow ($u_{\text{rev}}/u_{\infty}=13\%$ for PIV compared to $7\%$ for RANS, where $u_{\text{rev}}$ denotes the peak reverse flow of the mean field). The RANS mean field displays a thin separation bubble in the upstream part which is not visible in the PIV mean field. This latter observation may be attributed to the presence of spurious measurement data near the wall in the current dataset, caused by low signal-to-noise ratios due to strong light reflections. This experimental shortcoming does not negatively affect the SPOD analysis, since the breathing mode unfolds further away from the wall, as will be shown later. With the chosen RANS model using the $k$-$\omega$-SST turbulence model and adaption of the turbulence intensity value at the inlet of the domain, it was not possible to better approximate the crosswise extent of the bubble without significantly altering the separation and reattachment points. Similar challenges in matching RANS-predicted separation bubble behaviour with experimental observations have been reported for smooth-body separation, where accurate prediction remains difficult and continues to be an active topic of research (williams_Experimental_2020; gray_Experimental_2023; Klopsch2025). SarrasJfM2024 observed comparable discrepancies and showed that improved agreement could be achieved by applying adjoint-based data assimilation with flexibility in the turbulence model.
@@ 3.2 Global Linear Stability Analysis
To investigate modal amplification mechanisms in the flow, a global linear stability analysis (LSA) of the turbulent mean flow is performed. This approach examines the time-asymptotic dynamics of harmonic perturbations that evolve around the temporal mean state by performing an eigenvalue decomposition of the linearized Navier–Stokes operator, which yields growth rates and corresponding mode shapes. Linear stability analysis has become a common framework to study coherent flow dynamics (Sipp2010; Barkley2006) and has recently been applied to separation bubble flows (Cura2024; Cura2025; SarrasJfM2024).